OpenStudy (anonymous):

Which is the smaller mass? 1 milligram 1 centigram 1 decigram 1 gram

OpenStudy (anonymous):

1 milligram

OpenStudy (anonymous):

1 milligram

OpenStudy (anonymous):

would anyone care to explain why?

OpenStudy (anonymous):

Prefix --------- gram - 1g ============== Suffix(s) --------- milli - 0.001 kilo - 1000 deci - 0.1 centi - 0.01

OpenStudy (anonymous):

1mg=.001gm 1cg=.01gm 1dg=.1gm

OpenStudy (anonymous):

Thanks much better:))

OpenStudy (anonymous):

1mg_.001gm 1cg=.01gm 1dg=.1gm

OpenStudy (anonymous):

because 1mg=1/1000g 1cg=1/100 1dg=1/10 so least value is 1mg

OpenStudy (anonymous):

1 mg smallest = example my capsule of vitamins has a 250mg each

OpenStudy (anonymous):

1 Milligram would be the smallest for this answer.